# Finite Difference Solution Ansatz approach in Least-Squares Monte Carlo

### v7

@@ 2.1.4 Finite difference approach in PDE
As opposed to LSM, alternative numerical approach to solve the American-style option pricing is the finite difference (FD) method. The elegance of this method lies in the fact that it marches from maturity to present, which naturally fits into the requirement of backward recursive optimization in stopping time problems. We restrict the discussion within low-dimensional only, i.e., 1D, for maximum efficiency and accuracy. The most generic model reliable in replicating the market characteristics and solvable under 1D FD PDE is of the form of local volatility, under which the option price $V^{FD}(S,t)$ as function of underlying spot $S$ and time $t$ satisfies:
@@
for issuer’s exercise, respectively.
As we can see later, for a general model setting beyond 1D mentioned in Sec.2.2.4, Eqn.2.16 represents a 1D PDE after dimensional reduction with the coefficients $q_{FD}(S,t)$ and $\sigma_{FD}(S,t)$, corresponding to the effective single-asset continuous dividend yield and local volatility, respectively, to be specified case by case, e.g., in Sec.2.2.5 for multi-variate local volatility and single-asset stochastic volatility.
@@ 2.2.1 Limiting Case
Before proceeding, we introduce below proposition as a limiting case:
Proposition 1.
In a generalized LSM formulism, the continuation value estimator is exact if the exact continuation value is used as the only regressor.
The proof of Prop.1 is omitted as it is an elementary consequence of conditional expectation as $L^{2}$ projection that minimizes the objective function of LSM. This conclusion is free from model setting assumption.
For a general problem of interest, one could find an auxiliary variable with existing exact (or quasi-exact) solution as an ansatz to start with as one of the basis functions in the $L^{2}$ projection, with the expectation of dominating regression weight on this ansatz in the spirit of Prop.1. Inspired by the simplicity of FD solution for 1D PDE in Sec.2.1.4 with efficiency and quasi-exact accuracy, we would intuitively take the option price under 1D process as auxiliary variable, whose FD solution of continuation value $f_{k}^{\text{FD}}(x)$ is given by Eqn.2.17. From now on, we call this new method by incorporating $f_{k}^{\text{FD}}(x)$ into one of the LSM basis functions as FD Solution Ansatz based Least-Squares Monte Carlo (FD-LSM).
Mathematically, FD-LSM is defined by a modified least-squares continuation estimator with a new $L^{2}$ projection operator $P_{k}^{\text{FD-LSM}}$, as opposed to the classic one $P_{k}^{\text{LSM}}$(previously denoted by $P$ defined in Eqn.2.2) operating on $\vec{\phi}_{R}^{\text{LSM}}(x)$(previously denoted by $\vec{\phi}^{P}_{R}(x)$). The new basis functions of $P_{k}^{\text{FD-LSM}}$ is given by $\vec{\phi}_{R}^{\text{FD-LSM}}(x):=\vec{\phi}^{\text{FD}}(x)\cup\vec{\phi}_{R-1}^{\text{LSM}}(x)$ as a union of $\vec{\phi}^{\text{FD}}(x)$ and $\vec{\phi}_{R-1}^{\text{LSM}}(x)$ ($R$ is decremented by one to cater for the added new FD ansatz), where $\vec{\phi}^{\text{FD}}(x)=\{1,f_{k}^{\text{FD}}(x)\}$ [^4] We allow a constant term in the regressors to cater for mismatch of center of mass between independent and dependent variables. under which the associated projection operator $P_{k}^{\text{FD}}$ is also introduced for illustrative purpose later.
@@ 2.2.2 FD Solution: Control Variate
Below proposition is established for the illustration of fundamental improvement of $P_{k}^{\text{FD-LSM}}$ over $P_{k}^{\text{LSM}}$ from a theoretical point of view.
@@
For notational convenience on random variables, we suppress the spatial dependence and condense the notation by replacing arguments depending on time by subscripts and further replace $t_{k}$ with simply $k$, and thus $V_{k+1}:=V(\omega,\tau(t_{k+1}))$.
@@
First of all, let’s denote by $\hat{V}_{k+1}^{1D}$ the FD solvable auxiliary variable with the interpretation as the path option value under 1D process satisfying $e^{-r\Delta T}\operatorname{\mathbb{E}}[\hat{V}_{k+1}^{1D}|X=x]=f_{k}^{\text{FD}}(x)$. It is the equivalent counterpart of $V_{k+1}$ but under a hypothetical 1D simulation process instead.
By projecting the original option price into $\vec{\phi}^{\text{FD}}(x)$, we have
$P_{k}^{\text{FD}}[e^{-r\Delta T}V_{k+1}]=\beta_{k}^{\text{FD}}f_{k}^{\text{FD}}(x)+\alpha_{k}^{\text{FD}},$
where from simple linear regression, the optimal coefficients are given by
$\begin{split}\beta_{k}^{\text{FD}}&=\frac{\operatorname{\mathbb{E}}[\operatorname{\mathrm{Cov}}(V_{k+1},\hat{V}_{k+1}^{1D}|\operatorname{\mathcal{F}}_{t_{k}})|\operatorname{\mathcal{F}}_{0}]}{\operatorname{\mathbb{E}}[\operatorname{\mathrm{Var}}(\hat{V}_{k+1}^{1D}|\operatorname{\mathcal{F}}_{t_{k}})|\operatorname{\mathcal{F}}_{0}]},\\ \alpha_{k}^{\text{FD}}&=e^{-r\Delta T}(\operatorname{\mathbb{E}}(V_{k+1}|\operatorname{\mathcal{F}}_{0})-\beta_{k}^{\text{FD}}\operatorname{\mathbb{E}}(\hat{V}_{k+1}^{1D}|\operatorname{\mathcal{F}}_{0})),\\ \end{split}$ (2.18)
with $\operatorname{\mathrm{Var}}(\cdot|\operatorname{\mathcal{F}}_{t_{k}})$ and $\operatorname{\mathrm{Cov}}(\cdot,\cdot\cdot|\operatorname{\mathcal{F}}_{t_{k}})$ denoting the conditional variance and covariance operators, respectively.
@@
Applying the projection $P_{k}^{\text{FD-LSM}}$ to $e^{-r\Delta T}V_{k+1}$ gives
@@
$\begin{split}P_{k}^{\text{FD-LSM}}[e^{-r\Delta T}V_{k+1}]&=P_{k}^{\text{FD}}[e^{-r\Delta T}V_{k+1}]+P_{k}^{\text{LSM}}[e^{-r\Delta T}V_{k+1}-P_{k}^{\text{FD}}[e^{-r\Delta T}V_{k+1}]]\\ &=\beta_{k}^{\text{FD}}f_{k}^{\text{FD}}(x)+\alpha_{k}^{\text{FD}}+P_{k}^{\text{LSM}}[e^{-r\Delta T}V_{k+1}-\beta_{k}^{\text{FD}}f_{k}^{\text{FD}}(x)-\alpha_{k}^{\text{FD}}]\\ &=\beta_{k}^{\text{FD}}f_{k}^{\text{FD}}(x)+P_{k}^{\text{LSM}}[e^{-r\Delta T}V_{k+1}-\beta_{k}^{\text{FD}}f_{k}^{\text{FD}}(x)],\end{split}$ (2.19)
@@
where in the last line, the constant term $\alpha_{k}^{\text{FD}}$ has been absorbed into $P_{k}^{\text{LSM}}$.
@@
In the limit of $R\rightarrow\infty$, we have $P_{k}^{\text{LSM}}[e^{-r\Delta T}\hat{V}_{k+1}^{1D}]=f_{k}^{\text{FD}}(x)$ due to convergence in LSM. Thus we arrive at the desired result as a consequence of linearity property of $L^{2}$ projection:
$P_{k}^{\text{FD-LSM}}[e^{-r\Delta T}V_{k+1}]=\beta_{k}^{\text{FD}}f_{k}^{\text{FD}}(x)+P_{k}^{\text{LSM}}[e^{-r\Delta T}(V_{k+1}-\beta_{k}^{\text{FD}}\hat{V}_{k+1}^{1D})],$ (2.20)
@@
which claims that $P_{k}^{\text{FD-LSM}}$ is equivalent to a customized $P_{k}^{\text{LSM}}$ on a control-variated “kernel” with the help of the FD ansatz. ∎
@@
Based on Prop.2, proof of convergence on LSM from previous studies (Longstaff and Schwartz, (2001); Stentoft, (2001)) is equally applicable to FD-LSM. As $R\rightarrow\infty$, both $P_{k}^{\text{FD-LSM}}$ and $P_{k}^{\text{LSM}}$ converge to the exact continuation function and are thus unbiased estimators. However, the variance of projection “kernel” in the former is reduced tremendously in a control variated manner by observing
@@
$\operatorname{\mathbb{E}}[\operatorname{\mathrm{Var}}(V_{k+1}-\beta_{k}^{\text{FD}}\hat{V}_{k+1}^{1D}|\operatorname{\mathcal{F}}_{t_{k}})|\operatorname{\mathcal{F}}_{0}]=[1-\tilde{\rho}^{2}(V_{k+1},\hat{V}_{k+1}^{1D})]\cdot\operatorname{\mathbb{E}}[\operatorname{\mathrm{Var}}(V_{k+1}|\operatorname{\mathcal{F}}_{t_{k}})|\operatorname{\mathcal{F}}_{0}]$ (2.21)
@@
$\tilde{\rho}(V_{k+1},\hat{V}_{k+1}^{1D})=\frac{\operatorname{\mathbb{E}}[\operatorname{\mathrm{Cov}}(V_{k+1},\hat{V}_{k+1}^{1D}|\operatorname{\mathcal{F}}_{t_{k}})|\operatorname{\mathcal{F}}_{0}]}{\sqrt{\operatorname{\mathbb{E}}[\operatorname{\mathrm{Var}}(V_{k+1}|\operatorname{\mathcal{F}}_{t_{k}})|\operatorname{\mathcal{F}}_{0}]\cdot\operatorname{\mathbb{E}}[\operatorname{\mathrm{Var}}(\hat{V}_{k+1}^{1D}|\operatorname{\mathcal{F}}_{t_{k}})|\operatorname{\mathcal{F}}_{0}]}}$ (2.22)
defining the correlation between $V_{k+1}$ and $\hat{V}_{k+1}^{1D}$.
Considering Eqn.2.4 together with above, one can see that least-squares error (lower bound) in FD-LSM is also reduced by the same scaling factor as the variance:
$\begin{split}\inf_{\vec{b}\in\mathbb{R}^{R}}\mathcal{L}^{P^{\text{FD-LSM}}}(\vec{b};e^{-r\Delta T}V_{k+1})&=\inf_{\vec{b}\in\mathbb{R}^{R}}\mathcal{L}^{P^{\text{LSM}}}(\vec{b};e^{-r\Delta T}(V_{k+1}-\beta_{k}^{\text{FD}}\hat{V}_{k+1}^{1D}))\\ &=[1-\tilde{\rho}^{2}(V_{k+1},\hat{V}_{k+1}^{1D})]\cdot\inf_{\vec{b}\in\mathbb{R}^{R}}\mathcal{L}^{P^{\text{LSM}}}(\vec{b};e^{-r\Delta T}V_{k+1}).\end{split}$ (2.23)

### v8

@@ 2.2.4 Finite difference approach in PDE
As opposed to LSM, alternative numerical approach to solve the American-style option pricing is the finite difference (FD) method. The elegance of this method lies in the fact that it marches from maturity to present, which naturally fits into the requirement of backward recursive optimization in stopping time problems. We restrict the discussion within 1D for maximum efficiency and accuracy. The local volatility setting as describe in Sec.2.1 with $d=1$ is solvable under 1D FD PDE, under which the option price $V^{FD}(S,t)$ as function of underlying spot $S$ (omitting the subscript $i$ for the mono underlying case) and time $t$ satisfies:
@@
for issuer’s exercise, respectively. As we will see later, using the approach of Gyöngy, (1986), for a more general higher-dimensional setting, the marginal distribution of original process after dimensional reduction is mimicked by a 1D PDE in Eqn.2.20 with the coefficients $q_{FD}(S,t)$ and $\sigma_{FD}(S,t)$ (re-labeled from $q_{1}(S,t)$ and $\sigma_{1}(S,t)$ in Eqn.2.1), corresponding to the effective single-asset continuous dividend yield and local volatility, respectively. The details of such dimensional reduction will specified case by case, e.g., for multi-variate local volatility and single-asset stochastic volatility, respectively, in Sec.2.3.4.
@@ 2.3.1 Limiting Case
Before proceeding, we introduce below lemma as a limiting case:
Lemma 1.
In a generalized LSM formulism, the continuation value estimator is exact if the exact continuation value is used as one of the regressors.
The proof of Lemma 1 is omitted as it is an elementary consequence of conditional expectation as $L^{2}$ projection that minimizes the objective function of LSM. This conclusion is free from model setting assumption.
For a general problem of interest, one could find an auxiliary variable with existing exact (or quasi-exact) solution as an ansatz to start with as one of the basis functions in the $L^{2}$ projection, with the expectation of dominating regression weight on this ansatz in the spirit of Lemma 1. Inspired by the simplicity of FD solution for 1D PDE in Sec.2.2.4 with efficiency and quasi-exact accuracy, we would intuitively take the option price under 1D process as auxiliary variable, whose FD solution of continuation value $f_{k}^{\text{FD}}(x)$ is given by Eqn.2.21. From now on, we call this new method by incorporating $f_{k}^{\text{FD}}(x)$ into one of the LSM basis functions as FD Solution Ansatz based Least-Squares Monte Carlo (FD-LSM).
@@ 2.3.2 FD Solution: Control Variate
Mathematically, FD-LSM is defined by a modified least-squares continuation estimator with a new $L^{2}$ projection operator $P_{k}^{\text{FD-LSM}}$, as opposed to the classic one $P_{k}^{\text{LSM}}$(previously denoted by $P$ defined in Eqn.2.4) operating on $\vec{\phi}_{R}^{P^{\text{LSM}}}(x)$(previously denoted by $\vec{\phi}^{P}_{R}(x)$). The new basis functions of $P_{k}^{\text{FD-LSM}}$ is given by $\vec{\phi}_{R}^{P^{\text{FD-LSM}}}(x):=\vec{\phi}^{P^{\text{FD}}}(x)\cup\vec{\phi}_{R-1}^{P^{\text{LSM}}}(x)$ ($R$ is decremented by one to cater for the added new FD ansatz), where $\vec{\phi}^{P^{\text{FD}}}(x):=\{1,f_{k}^{\text{FD}}(x)\}^{T}$ [^4] We allow a constant term in the regressors to cater for mismatch of center of mass between independent and dependent variables. under which the associated projection operator $P_{k}^{\text{FD}}$ is also introduced for illustrative purpose later.
From the point on, for notational convenience on random variables, we suppress the spatial dependence and condense the notation by replacing arguments depending on time by subscripts and further replace $t_{k}$ with simply $k$ if applicable. For example, $X_{k}\equiv X(\omega,t_{k})$, $\vec{\phi}^{P}_{R,k}\equiv\vec{\phi}^{P}_{R}(X_{k})$, $\vec{\phi}^{P^{\text{FD}}}_{k}\equiv\vec{\phi}^{P^{\text{FD}}}(X_{k})$, $V_{k+1}\equiv V(\omega,\tau(t_{k+1}))$, $Z_{k}\equiv Z(\omega,t_{k})$, $F_{k}\equiv F(\omega,t_{k})$ and $F^{P}_{k}\equiv F^{P}(\omega,t_{k})$.
To formally define $P_{k}^{\text{FD}}$, let’s denote by $\hat{V}_{k+1}^{1D}$ the FD solvable auxiliary variable with the interpretation as the path option value under 1D process satisfying $e^{-r\Delta T}\operatorname{\mathbb{E}}[\hat{V}_{k+1}^{1D}|X=x]=f_{k}^{\text{FD}}(x)$. It is the equivalent counterpart of $V_{k+1}$ but under a hypothetical 1D simulation process instead.
At $t_{k}$, by projecting a given stochastic variable $Y_{t\geq t_{k}}$ onto $\text{span}(\vec{\phi}^{P^{\text{FD}}}_{k})$, defined similarly as Eqn.2.6, we have
$P_{k}^{\text{FD}}[Y_{t}]:=\beta_{k}^{\text{FD}}(Y_{t})f_{k}^{\text{FD}}(X_{k})+\alpha_{k}^{\text{FD}}(Y_{t}),$ (2.22)
where from simple linear regression, the optimal coefficients as functions of $Y_{t}$ are given by
$\begin{split}\beta_{k}^{\text{FD}}(Y_{t})&:=\frac{\operatorname{\mathbb{E}}[\operatorname{\mathrm{Cov}}(Y_{t},\hat{V}_{k+1}^{1D}|\operatorname{\mathcal{F}}_{t_{k}})|\operatorname{\mathcal{F}}_{0}]}{\operatorname{\mathbb{E}}[\operatorname{\mathrm{Var}}(\hat{V}_{k+1}^{1D}|\operatorname{\mathcal{F}}_{t_{k}})|\operatorname{\mathcal{F}}_{0}]},\\ \alpha_{k}^{\text{FD}}(Y_{t})&:=\operatorname{\mathbb{E}}(Y_{t}|\operatorname{\mathcal{F}}_{0})-\beta_{k}^{\text{FD}}(Y_{t})e^{-r\Delta T}\operatorname{\mathbb{E}}(\hat{V}_{k+1}^{1D}|\operatorname{\mathcal{F}}_{0}),\end{split}$ (2.23)
with $\operatorname{\mathrm{Var}}(\cdot|\operatorname{\mathcal{F}}_{t_{k}})$ and $\operatorname{\mathrm{Cov}}(\cdot,\cdot\cdot|\operatorname{\mathcal{F}}_{t_{k}})$ denoting the conditional variance and covariance operators, respectively. From Eqn.2.23, we can also see that
$P_{k}^{\text{FD}}[F_{k}]=\beta_{k}^{\text{FD}}(F_{k})f_{k}^{\text{FD}}(X_{k})+\alpha_{k}^{\text{FD}}(F_{k})=P_{k}^{\text{FD}}[e^{-r\Delta T}V_{k+1}]$ (2.24)
because $\beta_{k}^{\text{FD}}(F_{k})=\beta_{k}^{\text{FD}}(e^{-r\Delta T}V_{k+1})$ and $\alpha_{k}^{\text{FD}}(F_{k})=\alpha_{k}^{\text{FD}}(e^{-r\Delta T}V_{k+1})$ given the $\operatorname{\mathcal{F}}_{t_{k}}$ filtration.
We also define the annihilator under $\vec{\phi}^{P^{\text{FD}}}(x)$ as $M_{k}^{\text{FD}}:=1-P_{k}^{\text{FD}}$. Both of $P_{k}^{\text{FD}}$ and $M_{k}^{\text{FD}}$ are idempotent.
With above definitions in place, it is ready to state and prove below lemma.
Lemma 2.
$P_{k}^{\text{FD-LSM}}$ on the option price can be decomposed exactly into $P_{k}^{\text{FD}}$ (defined in above) and $P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}$: $P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}$ utilizes the same basis functions as $P_{k}^{\text{LSM}}$, but the regression coefficients are obtained as the “net” effect of $\vec{\phi}_{R-1}^{P^{\text{LSM}}}(x)$ on $V_{k+1}$ with the effect of $\vec{\phi}^{P^{\text{FD}}}(x)$ annihilated from the remaining system.
Proof.
Here we are trying to partition a $L^{2}$ projection into 2 sets of explanatory variables, $\vec{\phi}^{P^{\text{FD}}}(x)$ and $\vec{\phi}_{R-1}^{P^{\text{LSM}}}(x)$. Based on the so-called Frisch-Waugh-Lovell (FWL) theorem  [^5] This theorem is well-recognized in the field of Econometrics, as a powerful tool for understanding and simplifying regression analysis. Its proof can be found in Section 2.4 of Davidson et al., (2004)., instead of regressing on all variables as a whole, one can perform two subsequent regressions on individual set: while the regression on $\vec{\phi}^{P^{\text{FD}}}(x)$ is done as usual, the regression coefficients of $\vec{\phi}_{R-1}^{P^{\text{LSM}}}(x)$ can be obtained equivalently by first “partialling out” the effects of $\vec{\phi}^{P^{\text{FD}}}(x)$ from both $V_{k+1}$ and $\vec{\phi}_{R-1}^{P^{\text{LSM}}}(x)$, and then running the regression of the partialled-out dependent variable, denoted by $V^{\perp\text{FD}}_{k+1}$, on the partialled-out regressors, denoted by $\vec{\phi}^{P^{\text{LSM}\perp\text{FD}}}_{R-1}(x)$.
In mathematical terms, $P_{k}^{\text{FD-LSM}}$ can be decomposed into $P_{k}^{\text{FD}}$ and $P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}$, where the latter is defined as
$P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}V_{k+1}]:=\vec{\beta}^{P^{\text{LSM}\perp\text{FD}}}_{k}\cdot\vec{\phi}^{P^{\text{LSM}}}_{R-1,k}$ (2.25)
with the coefficients obtained by
$\begin{split}\vec{\beta}^{P^{\text{LSM}\perp\text{FD}}}_{k}&=\arg_{\vec{b}\in\mathbb{R}^{R-1}}\min\mathcal{L}(e^{-r\Delta T}V_{k+1}^{\perp\text{FD}};\vec{b};\vec{\phi}^{P^{\text{LSM}\perp\text{FD}}}_{R-1,k}).\end{split}$ (2.26)
Here $\mathcal{L}$ adopts the same functional form as Eqn.2.5 but operates on below arguments by annihilating the effect of $\vec{\phi}^{P^{\text{FD}}}(x)$
$\begin{split}e^{-r\Delta T}V_{k+1}^{\perp\text{FD}}:&=M_{k}^{\text{FD}}[e^{-r\Delta T}V_{k+1}],\\ \vec{\phi}^{P^{\text{LSM}\perp\text{FD}}}_{R-1,k}:&=\{M_{k}^{\text{FD}}[\phi^{P^{\text{LSM}}}_{r,k}]|r\in[1,R-1]\}^{T}=M_{k}^{\text{FD}}[\vec{\phi}^{P^{\text{LSM}}}_{R-1,k}].\end{split}$
Similar to Eqn.2.8, the formal solution of $\vec{\beta}^{P^{\text{LSM}\perp\text{FD}}}_{k}$ as a function of $e^{-r\Delta T}V_{k+1}$ is
$\begin{split}\vec{\beta}^{P^{\text{LSM}\perp\text{FD}}}_{k}(e^{-r\Delta T}V_{k+1}):=(\operatorname{\mathbb{E}}[\vec{\phi}^{P^{\text{LSM}\perp\text{FD}}}_{R-1,k}\vec{\phi}^{P^{\text{LSM}\perp\text{FD}}T}_{R-1,k}|\operatorname{\mathcal{F}}_{0}])^{-1}\operatorname{\mathbb{E}}[\vec{\phi}^{P^{\text{LSM}\perp\text{FD}}}_{R-1,k}e^{-r\Delta T}V_{k+1}^{\perp\text{FD}}|\operatorname{\mathcal{F}}_{0}].\end{split}$ (2.27)
Due to the idempotent property of $M_{k}^{\text{FD}}$, we have $\vec{\beta}^{P^{\text{LSM}\perp\text{FD}}}_{k}(e^{-r\Delta T}V_{k+1})=\vec{\beta}^{P^{\text{LSM}\perp\text{FD}}}_{k}(e^{-r\Delta T}V_{k+1}^{\perp\text{FD}})$ and $P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}V_{k+1}]=P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}V^{\perp\text{FD}}_{k+1}]$. Thus, we arrive at
$\begin{split}P_{k}^{\text{FD-LSM}}[e^{-r\Delta T}V_{k+1}]&=P_{k}^{\text{FD}}[e^{-r\Delta T}V_{k+1}]+P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}V_{k+1}]\\ &=P_{k}^{\text{FD}}[e^{-r\Delta T}V_{k+1}]+P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}V_{k+1}^{\perp\text{FD}}].\end{split}$ (2.28)
@@
With the help of above lemma, below proposition can be established for the illustration of fundamental improvement of $P_{k}^{\text{FD-LSM}}$ over $P_{k}^{\text{LSM}}$ from a theoretical point of view.
@@
First of all, based on Lemma 2, let’s apply FD-LSM to the path option value under 1D process, which gives,
$\begin{split}P_{k}^{\text{FD-LSM}}[e^{-r\Delta T}\hat{V}_{k+1}^{1D}]&=P_{k}^{\text{FD}}[e^{-r\Delta T}\hat{V}_{k+1}^{1D}]+P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}\hat{V}_{k+1}^{1D}-P_{k}^{\text{FD}}[e^{-r\Delta T}\hat{V}_{k+1}^{1D}]]\\ &=f_{k}^{\text{FD}}(X_{k})+P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}\hat{V}_{k+1}^{1D}-f_{k}^{\text{FD}}(X_{k})].\end{split}$ (2.29)
Based on Lemma 1, the left-hand side in above is equal to $f_{k}^{\text{FD}}(X_{k})$, which implies
$P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}\hat{V}_{k+1}^{1D}-f_{k}^{\text{FD}}(X_{k})]=0.$ (2.30)
@@
Next, applying the projection $P_{k}^{\text{FD-LSM}}$ to the original option price $e^{-r\Delta T}V_{k+1}$ gives
@@
$\begin{split}P_{k}^{\text{FD-LSM}}[e^{-r\Delta T}V_{k+1}]&=P_{k}^{\text{FD}}[e^{-r\Delta T}V_{k+1}]+P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}V_{k+1}-P_{k}^{\text{FD}}[e^{-r\Delta T}V_{k+1}]]\\ &=\beta_{k}^{\text{FD}}(e^{-r\Delta T}V_{k+1})f_{k}^{\text{FD}}(X_{k})+\alpha_{k}^{\text{FD}}(e^{-r\Delta T}V_{k+1})\\ &\quad+P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}V_{k+1}-\beta_{k}^{\text{FD}}(e^{-r\Delta T}V_{k+1})f_{k}^{\text{FD}}(X_{k})-\alpha_{k}^{\text{FD}}(e^{-r\Delta T}V_{k+1})]\\ &=\beta_{k}^{\text{FD}}(F_{k})f_{k}^{\text{FD}}(X_{k})+P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}V_{k+1}-\beta_{k}^{\text{FD}}(F_{k})f_{k}^{\text{FD}}(X_{k})],\end{split}$ (2.31)
@@
where in the last line, $\beta_{k}^{\text{FD}}(e^{-r\Delta T}V_{k+1})$ has been replaced with $\beta_{k}^{\text{FD}}(F_{k})$ equivalently from Eqn.2.23 that leads to Eqn.2.24, and constant term $\alpha_{k}^{\text{FD}}(F_{k})$ has been absorbed into $P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}$.
@@
Given Eqn.2.30, we have $P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}\hat{V}_{k+1}^{1D}]=P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[f_{k}^{\text{FD}}(X_{k})]$. Thus we arrive at the desired result as a consequence of linearity property of $L^{2}$ projection:
$P_{k}^{\text{FD-LSM}}[e^{-r\Delta T}V_{k+1}]=\beta_{k}^{\text{FD}}(F_{k})f_{k}^{\text{FD}}(X_{k})+P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}(V_{k+1}-\beta_{k}^{\text{FD}}(F_{k})\hat{V}_{k+1}^{1D})],$ (2.32)
@@
which claims that $P_{k}^{\text{FD-LSM}}$ is equivalent to a customized projection on a control-variated “kernel” with the help of the FD ansatz. ∎
@@
Based on Prop.1, proof of convergence on LSM from previous studies (Longstaff and Schwartz, (2001); Stentoft, (2001)) is equally applicable to FD-LSM. Both $P_{k}^{\text{FD-LSM}}$ and $P_{k}^{\text{LSM}}$ are unbiased estimators. However, the variance of projection “kernel” in the former is reduced tremendously in a control variated manner by observing
@@
$\operatorname{\mathbb{E}}[\operatorname{\mathrm{Var}}(V_{k+1}-\beta_{k}^{\text{FD}}(F_{k})\hat{V}_{k+1}^{1D}|\operatorname{\mathcal{F}}_{t_{k}})|\operatorname{\mathcal{F}}_{0}]=[1-\tilde{\rho}_{k}^{2}(V_{k+1},\hat{V}_{k+1}^{1D})]\cdot\operatorname{\mathbb{E}}[\operatorname{\mathrm{Var}}(V_{k+1}|\operatorname{\mathcal{F}}_{t_{k}})|\operatorname{\mathcal{F}}_{0}]$ (2.33)
@@
$\tilde{\rho}_{k}(V_{k+1},\hat{V}_{k+1}^{1D}):=\frac{\operatorname{\mathbb{E}}[\operatorname{\mathrm{Cov}}(V_{k+1},\hat{V}_{k+1}^{1D}|\operatorname{\mathcal{F}}_{t_{k}})|\operatorname{\mathcal{F}}_{0}]}{\sqrt{\operatorname{\mathbb{E}}[\operatorname{\mathrm{Var}}(V_{k+1}|\operatorname{\mathcal{F}}_{t_{k}})|\operatorname{\mathcal{F}}_{0}]\cdot\operatorname{\mathbb{E}}[\operatorname{\mathrm{Var}}(\hat{V}_{k+1}^{1D}|\operatorname{\mathcal{F}}_{t_{k}})|\operatorname{\mathcal{F}}_{0}]}}$ (2.34)
defining the correlation between $V_{k+1}$ and $\hat{V}_{k+1}^{1D}$ at time step $t_{k}$.
Considering Eqn.2.7 (re-labeling $P$ to $P^{\text{LSM}}$) and Eqn.2.26 together with above, one can see that the RSS (lower bound) in FD-LSM is also reduced by the same scaling factor as the variance:
$\begin{split}\inf_{\vec{b}\in\mathbb{R}^{R}}\mathcal{L}(e^{-r\Delta T}V_{k+1};\vec{b};\vec{\phi}^{P^{\text{FD-LSM}}}_{R,k})&=\inf_{\vec{b}\in\mathbb{R}^{R-1}}\mathcal{L}(e^{-r\Delta T}(V_{k+1}-\beta_{k}^{\text{FD}}(F_{k})\hat{V}_{k+1}^{1D});\vec{b};\vec{\phi}^{P^{\text{LSM}\perp\text{FD}}}_{R-1,k})\\ &=\operatorname{\mathbb{E}}[\operatorname{\mathrm{Var}}(e^{-r\Delta T}V_{k+1}-e^{-r\Delta T}\beta_{k}^{\text{FD}}(F_{k})\hat{V}_{k+1}^{1D}|\operatorname{\mathcal{F}}_{t_{k}})|\operatorname{\mathcal{F}}_{0}]\\ &=[1-\tilde{\rho}_{k}^{2}(V_{k+1},\hat{V}_{k+1}^{1D})]\cdot\operatorname{\mathbb{E}}[\operatorname{\mathrm{Var}}(e^{-r\Delta T}V_{k+1}|\operatorname{\mathcal{F}}_{t_{k}})|\operatorname{\mathcal{F}}_{0}]\\ &=[1-\tilde{\rho}_{k}^{2}(V_{k+1},\hat{V}_{k+1}^{1D})]\cdot\inf_{\vec{b}\in\mathbb{R}^{R}}\mathcal{L}(e^{-r\Delta T}V_{k+1};\vec{b};\vec{\phi}^{P^{\text{LSM}}}_{R,k}).\end{split}$ (2.35)